# Beyond Trial-and-Error: Predicting User Abandonment After a Moderation Intervention

### v2

@@ VII-B7 Model training
VII-C Model evaluation
We evaluate the optimized models and the baselines on the test set by means of standard evaluation metrics for classification tasks. Specifically, we report averaged scores such as the micro and macro F1 scores. Additionally, we report the precision, recall, and F1 score of the positive class (i.e, abandoning users), given its importance for moderators and platform administrators in the context of planning moderation interventions.
@@ VII-C Model evaluation
task model p precision recall F1 macro F1 micro F1
@@
Stratified 0.155 0.158 0.157 0.505 0.751
DT Ratio 1 0.356 0.199 0.255 0.530 0.689
DT Trend 1 0.325 0.219 0.262 0.547 0.726
@@
Stratified 0.263 0.266 0.264 0.497 0.605
DT Ratio 1 0.374 0.328 0.349 0.543 0.624
DT Trend 1 0.282 0.319 0.299 0.530 0.644
@@
TABLE II: Classification results for the hard abandonment (top rows) and soft abandonment (bottom rows) tasks. Column p reports the number of features used by each model. For each task, the best result in each evaluation metric is shown in bold and the second-best is underlined.
VIII Results
VIII-A Prediction of hard and soft abandonment
@@
Classification results for the hard abandonment task are reported in the top half of Table II, while those for the soft abandonment task are reported in the bottom half of the table. For each task we report the performance of the baselines and of the optimized version of each trained model. For each baseline and model, Table II shows the number of used features (p) and the evaluation metrics computed for the positive class and for both classes (overall columns). With respect to the hard abandonment task, Gradient Boosting (GB) achieved the best results all-round. It was the best performing model in each evaluation metric, with the exception of precision on the positive class, where it was surpassed by Decision Tree (DT) and Adaptive Boost (AB). The second-best model in this task is Support Vector Machine (SVM) that achieved competitive results in all metrics and second-best results in positive class recall, overall macro F1, and overall micro F1. Despite the relatively solid performance exhibited by some models (e.g., GB and SVM), the results from Table II demonstrate the difficulty of this task. First of all, in general the results obtained by the trained models were encouraging but not exceptionally good, with the best result being micro F1 $=0.800$ by GB. Second, the difficulty faced by the trained models is also testified by the small margin by which some of the models —such as K-Nearest Neighbors (KNN) and Naive Bayes (NB)— outperformed the baselines. Finally, the results in table highlight a clear trend where the more complex models consistently outperformed the simpler ones. This shows that the increased complexity of models such as GB, SVM, and AB was needed in order to obtain more accurate predictions. However, in spite of the difficulty of the task, the majority of optimized models used a relatively low number of features. Indeed, with the exception of SVM and RF, all other models used $p\leq 20$ features, out of the 142 total ones that we computed. This result implies that many features conveyed little or redundant information for the task. The scores reported in the bottom half of Table II reveal more nuanced results for the soft abandonment task. A first interesting finding is that the overall results are worse for the soft abandonment task than for hard abandonment, while those for the positive class are better. For example, the best micro F1 $=0.737$ in soft abandonment, versus best micro F1 $=0.800$ in hard abandonment. Similarly, the best macro F1 $=0.664$ in soft abandonment, versus best macro F1 $=0.676$ in hard abandonment. On the contrary, the best positive class F1 $=0.516$ in soft abandonment, which is higher than F1 $=0.474$ in hard abandonment. In other words, by moving from hard to soft abandonment we improved predictions on the positive class (i.e., the abandoning users) but we degraded those on the negative one (i.e., the users who remained active). There are two related reasons for this behavior: (i) the definition of abandoning users and (ii) class imbalance. For the former, we recall that the soft abandonment task was introduced precisely to address the issue with those users who maintain some activity in the aftermath of the ban, but that eventually leave the platform, which represent challenging instances to classify. Therefore, their labeling in the soft abandonment task as abandoning users (positive class) rather than active ones (negative class), led to a certain degree of improvement of the models on the positive class. The second related factor is about class imbalance given that, before rebalancing, the soft abandonment task is less imbalanced than hard abandonment: 73/27 versus 85/15, respectively. To this end, we recall that abandoning users are the positive and the minority class in both tasks. As such, models trained for detecting soft abandonment train on a relatively larger number of positive instances than those for hard abandonment, hence the better performance on the positive class. Besides, results for the soft abandonment task are nuanced also in terms of the best performing models. In fact, there is no clear winner in this task, with SVM and GB —and to a lower extent, RF— achieving comparable performance. Nonetheless, the trend about the better performance of the more complex models is confirmed also in the soft abandonment task.
VIII-B Predicting abandonment at different levels of user activity
@@
task level of activity pos. neg. positive F1 macro F1 micro F1
hard aban. VH 731 $<$ n 353 2,896 0.400 0.656 0.848
HI 334 $<$ n $\leq$ 731 386 2,858 0.402 0.650 0.826
ME 156 $<$ n $\leq$ 334 432 2,801 0.446 0.665 0.808
LO 55 $<$ n $\leq$ 156 464 2,783 0.460 0.666 0.794
VL n $\leq$ 55 793 2,477 0.517 0.643 0.688
soft aban. VH 731 $<$ n 764 2,485 0.487 0.666 0.736
HI 334 $<$ n $\leq$ 731 742 2,502 0.519 0.691 0.788
ME 156 $<$ n $\leq$ 334 759 2,474 0.577 0.725 0.805
LO 55 $<$ n $\leq$ 156 833 2,414 0.521 0.676 0.751
VL n $\leq$ 55 1,286 1,984 0.600 0.639 0.644
@@
TABLE III: Classification results of Gradient Boosting for the hard abandonment (top rows) and soft abandonment (bottom rows) tasks at different levels of activity of the users. For each level of activity we report the corresponding range of comments n, the number of users in each class, and the evaluation metrics: F1 score of the positive class, overall macro F1, overall micro F1. For each task, the best result in each evaluation metric is shown in bold and the second-best is underlined.
@@
To perform this analysis, we examine the distribution of user comments pre-ban and we assign each user to one of the following five activity levels: very low (VL), low (LO), medium (ME), high (HI), and very high (VH) activity. The five activity levels are obtained by binning the distribution of user comments at regular intervals of 20 quantiles each, so that VL $\leq Q_{20}$, $Q_{20}<$ LO $\leq Q_{40}$, and so on up to VH $>Q_{80}$. We then perform a 80/20 train-test split of the dataset stratifying not only for class labels but also for the newly defined activity levels. Finally, we pick the best performing model based on the results in Table II —that is, Gradient Boosting (GB)— we train it on the training split, and we test it on the held-out data. The results of this analysis are reported in Table III for both the hard (top rows) and soft abandonment (bottom rows) task. In table are reported the number of comments n in each activity level, the corresponding number of abandoning (positive class) and remaining (negative class) users, and the evaluation metrics: the F1 score of the positive class, the overall macro F1, and the overall micro F1. The number of users in the positive and negative classes, reported in Table III for each activity level and for each task, shows that the imbalance in the dataset increases when considering increasingly active users. As we already observed in the results of Table II, class imbalance in the data has a detrimental effect on the classification performance on the minority class (i.e., the positive class, corresponding to the abandoning users). In Table III, this is reflected by the fact that the highest positive F1 scores are obtained at the lowest levels of user activity, for both tasks. Indeed, the best positive F1 $=0.600$ is achieved on the soft abandonment task for users with very low (VL) activity, which is the experimental condition with the lowest class imbalance (61/39). In contrast, the best overall results are obtained for medium to high activity levels. The best all-round result is achieved in the hard abandonment task for users with very high (VH) activity, with micro F1 $=0.848$. Then, in the soft abandonment task the best results are obtained at the medium (ME) and high (HI) activity levels, with micro F1 $=0.805$ and $0.788$, respectively.
VIII-C Leave-one-out cross-validation
@@
task test subreddit users max overlap precision recall F1 macro F1 micro F1
hard abandonment CTH r/ChapoTrapHouse 8,471 0.2% 0.206 0.465 0.286 0.581 0.790
TD r/The_Donald 3,539 4.1% 0.481 0.508 0.494 0.666 0.755
CP r/ConsumeProduct 1,098 15.0% 0.506 0.650 0.569 0.686 0.729
DHM r/DarkHumorAndMemes 929 3.8% 0.202 0.527 0.292 0.544 0.683
GC r/GenderCritical 918 1.4% 0.570 0.160 0.250 0.524 0.681
DAR r/DebateAltRight 325 40.0% 0.536 0.300 0.385 0.588 0.687
SNS r/ShitNeoconsSay 226 75.0% 0.653 0.368 0.471 0.618 0.676
TNR r/TheNewRight 144 49.0% 0.350 0.500 0.412 0.613 0.718
SB r/soyboys 142 73.0% 0.559 0.452 0.500 0.659 0.732
CCJ2 r/CCJ2 110 4.5% 0.556 0.294 0.385 0.650 0.852
DJC r/DarkJokeCentral 94 24.0% 0.176 0.600 0.273 0.517 0.640
WJ r/Wojak 57 49.0% 0.500 0.154 0.235 0.548 0.764
OEF r/OandAExclusiveForum 37 5.4% 0.143 1.000 0.250 0.522 0.676
HCH r/HateCrimeHoaxes 22 32.0% 0.667 0.200 0.308 0.509 0.591
ITH2 r/imgoingtohellforthis2 10 30.0% 0.000 0.000 0.000 0.438 0.778
averaged LOOCV results 0.407 $\pm$ 0.202 0.412 $\pm$ 0.237 0.341 $\pm$ 0.137 0.577 $\pm$ 0.069 0.717 $\pm$ 0.063
training/testing on all subreddits (no LOOCV) 0.603 0.392 0.474 0.676 0.800
soft abandonment CTH r/ChapoTrapHouse 8,471 0.2% 0.236 0.809 0.365 0.464 0.481
TD r/The_Donald 3,539 4.1% 0.658 0.541 0.594 0.684 0.709
CP r/ConsumeProduct 1,098 15.0% 0.701 0.473 0.565 0.651 0.672
DHM r/DarkHumorAndMemes 929 3.8% 0.444 0.446 0.445 0.626 0.713
GC r/GenderCritical 918 1.4% 0.739 0.213 0.331 0.520 0.594
DAR r/DebateAltRight 325 40.0% 0.697 0.317 0.436 0.533 0.554
SNS r/ShitNeoconsSay 226 75.0% 0.738 0.477 0.579 0.594 0.595
TNR r/TheNewRight 144 49.0% 0.571 0.483 0.523 0.618 0.641
SB r/soyboys 142 73.0% 0.711 0.403 0.514 0.615 0.641
CCJ2 r/CCJ2 110 4.5% 0.438 0.212 0.286 0.538 0.676
DJC r/DarkJokeCentral 94 24.0% 0.297 0.733 0.423 0.592 0.663
WJ r/Wojak 57 49.0% 0.778 0.280 0.412 0.574 0.636
OEF r/OandAExclusiveForum 37 5.4% 0.455 0.714 0.556 0.706 0.784
HCH r/HateCrimeHoaxes 22 32.0% 1.000 0.250 0.400 0.545 0.591
ITH2 r/imgoingtohellforthis2 10 30.0% 0.000 0.000 0.000 0.357 0.556
averaged LOOCV results 0.564 $\pm$ 0.245 0.423 $\pm$ 0.213 0.429 $\pm$ 0.146 0.574 $\pm$ 0.085 0.633 $\pm$ 0.073
training/testing on all subreddits (no LOOCV) 0.538 0.496 0.516 0.664 0.728
@@
TABLE IV: Results of the leave-one-out cross-validation (LOOCV) analysis. The top half of the table reports results for the hard abandonment task, while the bottom half is related to the soft abandonment. For both tasks, each row shows the classification results obtained with the reported subreddit used as test set and the remaining ones as training set. For each task, the bottom rows report the LOOCV results aggregated over all subreddits (mean $\pm$ standard deviation) and the reference results obtained by the same model without the LOOCV.
@@
Table IV shows the results of the LOOCV for the hard (top rows) and soft abandonment (bottom rows) tasks. Each row in the table reports testing results on a specific subreddit, when training on all others. In addition to the evaluation metrics, for each subreddit we also report the number of participating users and its maximum overlap with the other subreddits. Finally, for each task, the bottom rows report evaluation results aggregated across all subreddits, in terms of mean and standard deviation of each metric, as well the scores obtained by the GB model trained and tested on data from all subreddits —that is, without the LOOCV— as reported in Table II. The latter scores serve as reference values with which to compare the LOOCV results to quantify the performance decrease to expect when classifying users from an unseen subreddit. The averaged results reported in Table IV for both tasks indicate a high variability in the evaluation metrics of the positive class, as demonstrated by the large standard deviation with respect to the means. For instance, in the hard abandonment task, the F1 score ranges from a minimum of $0.235$ for r/Wojak (WJ) to a maximum of $0.569$ for r/ConsumeProduct (CP). Similarly, the recall ranges from $0.160$ for r/GenderCritical (GC) to $1.000$ for r/OandAExclusiveForum (OAE). The same can be observed for the soft abandonment task as well, where the precision ranges from $0.236$ for r/ChapoTrapHouse (CTH) to $1.000$ for r/HateCrimeOaxes (HCH). These results are indicative of substantial differences among the banned subreddits. In fact, in addition to the overall limited user overlap observed in Figure 4, the results from Table IV highlight notable differences in the behavior exhibited by the participants of such subreddits. Interestingly enough, the maximum overlap that each subreddit has with others, has little influence on the performance of the trained model. For each task, we computed the Pearson correlation coefficient r to estimate the strength of the linear relationship between the max overlap and three evaluation metrics —namely, the F1 score of the positive class, the macro F1, and the micro F1. The largest value measured for the hard abandonment task is r $=0.237$ (p $=0.394$), while the largest measured for the soft abandonment task is r $=0.222$ (p $=0.493$), both of which indicate weak and non-significant correlations. In turn, this strengthens the results reported in Table IV, including those about subreddit diversity, as the variable results in the table cannot be simply explained by the fact that some users participated in multiple subreddits, but rather to their inherent differences. The comparison between the results obtained without LOOCV and the LOOCV aggregate results sheds light on the capacity of our model to generalize to unseen data. To this end, results in Table IV show a moderate loss in performance in the LOOCV experiment, which is expected given the aforementioned differences among the subreddits. However, the percentage loss is generally contained. For example, in the hard abandonment task the average F1 score on the positive class obtained with the LOOCV is 71.9% of the original one. The loss for the overall metrics is even lower, with the average macro F1 and micro F1 being respectively 85.3% and 89.6% of the original respective metrics. Similar percentage losses are obtained for the soft abandonment task in the overall metrics, with the LOOCV average macro F1 and micro F1 maintaining respectively 86.5% and 87.0% of the original performance. Overall, these figures suggest that the model maintains encouraging performance even when applied to unseen data.
(a) Hard abandonment.
(b) Soft abandonment.

### v3

@@ 7.2.7 Model training
7.3 Model evaluation
We evaluate the optimized models and the baselines on the test set by means of standard evaluation metrics for classification tasks. Specifically, we report the area under the receiver operating characteristic curve (AUC), the overall micro F1 score, as well as the precision, recall, and F1 score of the positive class (i.e, abandoning users), given its importance for moderators and platform administrators in the context of planning moderation interventions. We also report the 95% confidence intervals of the positive F1 and the overall micro F1 scores.
@@ 7.3 Model evaluation
task model p precision recall F1 95% CI AUC micro F1 95% CI
@@
Stratified 0.155 0.158 0.157 - 0.501 0.751 -
DT Ratio 1 0.356 0.199 0.255 - 0.537 0.689 -
DT Trend 1 0.325 0.219 0.262 - 0.591 0.726 -
NB-NFE 142 0.206 0.313 0.248 - 0.595 0.814 -
DT-NFE 142 0.636 0.230 0.338 - 0.630 0.627 -
@@
Stratified 0.263 0.266 0.264 - 0.509 0.605 -
DT Ratio 1 0.374 0.328 0.349 - 0.527 0.624 -
DT Trend 1 0.282 0.319 0.299 - 0.549 0.644 -
NB-NFE 142 0.091 0.540 0.156 - 0.604 0.734 -
DT-NFE 142 0.589 0.287 0.386 - 0.524 0.494 -
@@
Table 2: Classification results for the hard abandonment (top rows) and soft abandonment (bottom rows) tasks. Column p reports the number of features used by each model. Model performance is measured in terms of precision, recall and F1 on the positive class, and in terms of overall AUC and micro F1. For positive F1 and overall micro F1 we also report the 95% confidence intervals based on the scores obtained in the k-fold cross-validation. For each task, the best result in each evaluation metric is shown in bold and the second-best is underlined.
(a) Hard abandonment.
(b) Soft abandonment.
Figure 5: Precision-Recall curves for the Gradient Boosting (GB) model in predicting hard and soft abandonment. The curves illustrate the trade-off between precision and recall, allowing for an assessment of how well the model balances false positives and false negatives in both tasks.
8 Results
8.1 Prediction of hard and soft abandonment
@@
Classification results for the hard abandonment task are reported in the top half of Table 2, while those for the soft abandonment task are reported in the bottom half of the table. For each task we report the performance of the baselines and of the optimized version of each trained model. For each baseline and model, Table 2 shows the number of used features (p) and the evaluation metrics computed for the positive class and for both classes (overall columns). With respect to the hard abandonment task, Gradient Boosting (GB) achieved the best results all-round. The second-best model in this task in terms of positive F1 is Random Forest (RF) that achieved competitive results in all metrics and second-best results in positive class recall and overall micro F1. Despite the relatively solid performance exhibited by some models (e.g., GB and RF), the results from Table 2 demonstrate the difficulty of this task. First of all, in general the results obtained by the trained models were encouraging but not exceptionally good, with the best result being micro F1 $=0.914$ by GB. Second, the difficulty faced by the trained models is also testified by the small margin by which some of the models —such as K-Nearest Neighbors (KNN) and Naive Bayes (NB)— outperformed the baselines, with Naive Bayes (NB) even having lower precision and positive F1 with respect to those achieved by DT-NFE in the hard abandonment task. However, for certain models our pipeline proved highly beneficial, significantly surpassing baseline performance as in the case of Support Vector Machine (SVM). Finally, results in table highlight a clear trend where the more complex models consistently outperformed the simpler ones. This shows that the increased complexity of models such as GB, RF, and AB was needed in order to obtain more accurate predictions. The scores reported in the bottom half of Table 2 reveal more nuanced results for the soft abandonment task. Again, the best-performing model is Gradient Boosting (GB), achieving positive F1 $=0.590$. Figure 5 shows precision-recall curves for this model, for both the hard and soft abandonment tasks. In hard abandonment, precision initially rises sharply and then stabilizes at a relatively high level for a wide range of recall values, before gradually decreasing. Conversely, in soft abandonment, while the initial trend is similar, precision declines more steeply as recall increases. This indicates that as the model attempts to retrieve more positive instances, it misclassifies a greater number of negative instances, leading to a faster deterioration in precision.
A first interesting finding is that the overall results are worse for the soft abandonment task than for hard abandonment, while those for the positive class are better. For example, the best micro F1 $=0.814$ in soft abandonment, versus best micro F1 $=0.914$ in hard abandonment. On the contrary, for the simpler models that underwent our pipeline the best positive F1 $=0.499$ in soft abandonment, which is higher than F1 $=0.434$ in hard abandonment. In other words, by moving from hard to soft abandonment we improved predictions on the positive class (i.e., the abandoning users) but we degraded those on the negative one (i.e., the users who remained active). There are two related reasons for this behavior: (i) the definition of abandoning users and (ii) class imbalance. For the former, we recall that the soft abandonment task was introduced precisely to address the issue with those users who maintain some activity in the aftermath of the ban, but that eventually leave the platform, which represent challenging instances to classify. Therefore, their labeling in the soft abandonment task as abandoning users (positive class) rather than active ones (negative class), led to a certain degree of improvement of the models on the positive class. The second related factor is about class imbalance given that, before rebalancing, the soft abandonment task is less imbalanced than hard abandonment: 73/27 versus 85/15, respectively. To this end, we recall that abandoning users are the positive and the minority class in both tasks. As such, models trained for detecting soft abandonment train on a relatively larger number of positive instances than those for hard abandonment, hence the better performance on the positive class.
8.2 Predicting abandonment at different levels of user activity
@@
task level of activity pos. neg. positive F1 micro F1
hard aban. VH 731 $<$ n 353 2,896 0.335 0.591
HI 334 $<$ n $\leq$ 731 386 2,858 0.726 0.929
ME 156 $<$ n $\leq$ 334 432 2,801 0.743 0.941
LO 55 $<$ n $\leq$ 156 464 2,783 0.263 0.871
VL n $\leq$ 55 793 2,477 0.548 0.768
soft aban. VH 731 $<$ n 764 2,485 0.493 0.813
HI 334 $<$ n $\leq$ 731 742 2,502 0.448 0.481
ME 156 $<$ n $\leq$ 334 759 2,474 0.536 0.842
LO 55 $<$ n $\leq$ 156 833 2,414 0.579 0.794
VL n $\leq$ 55 1,286 1,984 0.599 0.685
@@
Table 3: Gradient Boosting (GB) classification results in the hard abandonment (top rows) and soft abandonment (bottom rows) tasks, for users with different activity levels. For each activity level, we train a different model and report the corresponding range of comments n, the number of users in each class, and the evaluation metrics: F1 score of the positive class, overall micro F1. For each task, the best result in each evaluation metric is shown in bold and the second-best is underlined.
@@
To perform this analysis, we examined the distribution of user comments pre-ban and we assign each user to one of the following five activity levels: very low (VL), low (LO), medium (ME), high (HI), and very high (VH) activity. The five levels are obtained by binning the distribution of user comments at regular intervals of 20 quantiles each, so that VL $\leq Q_{20}$, $Q_{20}<$ LO $\leq Q_{40}$, and so on up to VH $>Q_{80}$. We then performed a 80/20 train-test split of the dataset stratifying not only for class labels but also for the newly defined activity levels. Finally, we picked the best performing model based on the results in Table 2 —that is, Gradient Boosting (GB)— and we trained and optimized a different model for each activity level. The results of this analysis are reported in Table 3 for both the hard (top rows) and soft abandonment (bottom rows) task. In table are reported the number of comments n in each activity level, the corresponding number of abandoning (positive class) and remaining (negative class) users, and the evaluation metrics: the F1 score of the positive class, and the overall micro F1. The number of users in the positive and negative classes, reported in Table 3 for each activity level and for each task, shows that the imbalance in the dataset increases when considering increasingly active users. As we already observed in the results of Table 2, class imbalance in the data has a detrimental effect on the classification performance on the minority class (i.e., the positive class, corresponding to the abandoning users). In Table 3, this is reflected by the fact that the best positive F1 $=0.599$ is achieved on the soft abandonment task for users with very low (VL) activity, which is the experimental condition with the lowest class imbalance (61/39). However, this is not true for the hard abandonment task, in which the lowest positive F1 is obtained for those users whose activity level is low (LO), while the best are obtained for high (HI) and medium (ME) activity levels. The best overall result is achieved in the hard abandonment task for users with medium (ME) activity, with micro F1 $=0.941$. Then, in the soft abandonment task the best overall results are obtained at the medium (ME) and very high (VH) activity levels, with micro F1 $=0.842$ and $0.813$, respectively.
Kendall’s $\tau$ Spearman’s $\rho$ RBO
task comparisons $p=10$ $p=\text{all}$ $p=10$ $p=\text{all}$ $p=10$ $p=\text{all}$
hard abandon. VH HI --0.241 0.231 --0.311 0.580 0.280 0.420
HI ME --0.009 0.022 --0.329 0.323 0.258 0.389
ME LO --0.222 0.018 --0.263 0.296 0.245 0.375
LO VL --0.205 0.128 --0.415 0.239 0.235 0.340
VH VL --0.205 0.231 --0.291 0.135 0.250 0.295
HI LO --0.402 0.005 --0.329 --0.272 0.258 0.310
soft abandon. VH HI --0.352 0.128 --0.178 0.191 0.239 0.389
HI ME --0.188 0.005 --0.329 --0.014 0.267 0.357
ME LO --0.365 0.018 --0.263 --0.013 0.267 0.342
LO VL --0.120 0.128 --0.415 --0.038 0.235 0.330
VH VL --0.239 0.231 --0.291 0.191 0.250 0.307
HI LO --0.402 0.005 --0.329 --0.272 0.258 0.310
$\dagger$ VH: Very High; HI: High; ME: Medium; LO: Low; VL: Very Low
Table 4: Agreement between the feature rankings of different models, each corresponding to users with a specific activity level. The comparisons column indicates which models are compared. Agreement between feature rankings is measured in terms of Kendall’s $\tau$ and Spearman’s $\rho$ rank correlations, and via the Rank-Biased Overlap (RBO) ranking similarity. Feature rankings are compared by considering all features ($p=\text{all}$) or only the top-10 most important ones ($p=10$).
After training a different model for each activity level, we assessed whether these models assigned similar importance to the features they used. We operationalized this analysis by computing several indicators of correlation and similarity between the feature rankings of the different models, such as Kendall’s $\tau$, Spearman’s $\rho$ and Rank-Biased Overlap (RBO) [72]. To rank the features of a model, we first identified and discarded strongly correlated features. In detail, we computed Pearson correlations r between each feature pair of a model, we identified those feature pairs for which $\textit{r}>|0.7|$, and we discarded one of such features at random. Then, for each model, we computed the feature importance of the remaining (uncorrelated) features with Permutation Importance (PI). PI measures the contribution of each feature to the model’s predictions by quantifying the performance degradation observed when its values are randomly shuffled while keeping all other features unchanged. Results are presented in Table 4 for the top-10 features of each model ($p=10$) and for all features ($p=\text{all}$). The column comparisons indicates which models —each corresponding to a different user activity level— are compared. As reported, we compared the feature ranking of models corresponding to similar activity levels (e.g., VH vs. HI), as well as those corresponding to opposite ones (e.g., VH vs. VL and HI vs. LO). The underlying intuition is that similar rankings should emerge among similar activity levels, whereas opposite levels are expected to exhibit less agreement.
For the hard abandonment task, results in Table 4 show that considering all features led to higher agreement, suggesting that the correlation tended to increase towards the tail of the rankings rather than the head. For instance, the correlation between VH and HI showed strong agreement, with Spearman’s $\rho=0.580$ and RBO $=0.420$ when $p=\text{all}$, but decreased when $p=10$. This suggests that highly active users share similar characteristics and can be grouped together. For the adjacent levels, this agreement decreases, particularly when transitioning from HI to ME and further down the activity scale. For instance, HI and ME showed a noticeable drop in agreement, with Kendall’s $\tau=-0.009$, indicating that the ranking of features differs significantly. This suggests that while highly active users (VH and HI) may exhibit similar behaviors, comparing similar but lower ranges of activity levels introduces more variability. The agreement between adjacent levels, such as HI and ME or ME and LO, was notably weaker, especially in Kendall’s $\tau$, which remains close to zero or even negative. Surprisingly, the agreement between VH and VL is stronger than some adjacent-level comparisons, with Spearman’s $\rho=0.135$ and RBO $=0.295$. While this correlation is still moderate, it suggests that these two groups, despite their opposing activity levels, might share some behavioral patterns. A possible explanation is that both represent users with extreme behaviors, leading to similar feature relevance.
Results in the soft abandonment task present the same patterns of the hard abandonment task. Higher agreement was generally found when all features are considered, with stronger correlations towards the tail of the rankings. For example, VH and HI exhibited relatively high agreement, particularly with respect to RBO, while Kendall’s $\tau$ remained lower, again highlighting sensitivity to rank shifts. Conversely, adjacent levels such as HI and ME, or ME and LO, show weak or even negative rank correlations, reinforcing the idea that mid-range activity users display more behavioral variability. Again, VH and VL show a correlation pattern similar to that of the hard abandonment task, with Spearman’s $\rho=0.191$ and RBO $=0.307$. This further supports the idea that extreme user groups, despite their contrasting activity levels, might share unexpected similarities. Overall, while some user groups exhibit similar feature importance rankings, others differ significantly. This highlights the need for tailored approaches rather than one-size-fits-all solutions. Understanding which users can be moderated similarly and which require separate treatment is pivotal for developing more personalized and effective moderation strategies [19].
8.3 Leave-one-out cross-validation
@@
task test subreddit users max overlap precision recall F1 AUC micro F1
hard abandonment CTH r/ChapoTrapHouse 8,471 0.2% .111 .859 .197 .685 .367
TD r/The_Donald 3,539 4.1% .801 .514 .626 .892 .855
CP r/ConsumeProduct 1,098 15.0% .806 .083 .151 .720 .741
DHM r/DarkHumorAndMemes 929 3.8% .737 .625 .676 .931 .926
GC r/GenderCritical 918 1.4% .548 .687 .610 .764 .708
DAR r/DebateAltRight 325 40.0% .769 .100 .177 .615 .697
SNS r/ShitNeoconsSay 226 75.0% .556 .851 .673 .808 .676
TNR r/TheNewRight 144 49.0% .000 .000 .000 .544 .789
SB r/soyboys 142 73.0% .800 .190 .308 .738 .746
CCJ2 r/CCJ2 110 4.5% .600 .882 .714 .961 .889
DJC r/DarkJokeCentral 94 24.0% 1.000 .100 .182 .587 .899
WJ r/Wojak 57 49.0% .846 .846 .846 .963 .927
OEF r/OandAExclusiveForum 37 5.4% .286 1.000 .444 .914 .865
HCH r/HateCrimeHoaxes 22 32.0% .714 .500 .588 .767 .682
ITH2 r/imgoingtohellforthis2 10 30.0% .000 .000 .000 .875 .889
averaged LOOCV results .571 $\pm$ .311 .482 $\pm$ .356 .412 $\pm$ .271 .784 $\pm$ .132 .777 $\pm$.141
training/testing on all subreddits (no LOOCV) .658 .736 .695 .930 .914
soft abandonment CTH r/ChapoTrapHouse 8,471 0.2% .184 .892 .305 .551 .250
TD r/The_Donald 3,539 4.1% .454 .712 .554 .608 .550
CP r/ConsumeProduct 1,098 15.0% .470 .582 .520 .517 .516
DHM r/DarkHumorAndMemes 929 3.8% .367 .648 .469 .680 .621
GC r/GenderCritical 918 1.4% .627 .747 .682 .744 .672
DAR r/DebateAltRight 325 40.0% .599 .617 .608 .553 .567
SNS r/ShitNeoconsSay 226 75.0% .598 .869 .708 .570 .581
TNR r/TheNewRight 144 49.0% .443 .879 .590 .525 .500
SB r/soyboys 142 73.0% .496 .896 .638 .536 .521
CCJ2 r/CCJ2 110 4.5% .346 .848 .491 .528 .463
DJC r/DarkJokeCentral 94 24.0% .204 .733 .319 .607 .472
WJ r/Wojak 57 49.0% .595 .880 .710 .624 .673
OEF r/OandAExclusiveForum 37 5.4% .217 .714 .333 .695 .459
HCH r/HateCrimeHoaxes 22 32.0% .556 .833 .667 .425 .545
ITH2 r/imgoingtohellforthis2 10 30.0% .333 .667 .444 .500 .444
averaged LOOCV results .432 $\pm$ .146 .767 $\pm$ .105 .535 $\pm$.135 .535 $\pm$.135 .522 $\pm$ .100
training/testing on all subreddits (no LOOCV) .513 .694 .590 .816 .808
@@
Table 5: Results of the leave-one-out cross-validation (LOOCV) analysis. The top half of the table reports results for the hard abandonment task, while the bottom half is related to the soft abandonment. For both tasks, each row shows the classification results obtained with the reported subreddit used as test set and the remaining ones as training set. For each task, the bottom rows report the LOOCV results aggregated over all subreddits (mean $\pm$ standard deviation) and the reference results obtained by the same model without the LOOCV.
@@
Table 5 shows the results of the LOOCV for the hard (top rows) and soft abandonment (bottom rows) tasks. Each row in the table reports testing results on a specific subreddit, when training on all others. In addition to the evaluation metrics, for each subreddit we also report the number of participating users and its maximum overlap with the other subreddits. Finally, for each task, the bottom rows report evaluation results aggregated across all subreddits, in terms of mean and standard deviation of each metric, as well the scores obtained by the GB model trained and tested on data from all subreddits —that is, without the LOOCV— as reported in Table 2. The latter scores serve as reference values with which to compare the LOOCV results to quantify the performance decrease to expect when classifying users from an unseen subreddit. The averaged results reported in Table 5 for both tasks indicate a high variability in the evaluation metrics of the positive class, as demonstrated by the large standard deviation with respect to the means. For instance, in the hard abandonment task, the F1 score ranges from a minimum of $0.000$ for r/TheNewRight (TNR) and r/imgoingtohellforthis2 (ITH2) to a maximum of $0.846$ for r/Wojak (WJ). Similarly, the recall ranges from $0.000$ for r/TheNewRight (TNR) and r/imgoingtohellforthis2 (ITH2) to $1.000$ for r/OandAExclusiveForum (OAE). The same can be observed for the soft abandonment task as well, where the precision ranges from $0.184$ for r/ChapoTrapHouse (CTH) to $0.827$ for r/GenderCritical (GC). These results are indicative of substantial differences among the banned subreddits. In fact, in addition to the overall limited user overlap observed in Figure 6, the results from Table 5 highlight notable differences in the behavior exhibited by the participants of such subreddits. Interestingly enough, the maximum overlap that each subreddit has with others, has little influence on the performance of the trained model. For each task, we computed the Pearson correlation coefficient r to estimate the strength of the linear relationship between the max overlap and two evaluation metrics —namely, the F1 score of the positive class and the micro F1. The largest value measured for the hard abandonment task is r $=0.237$ (p $=0.394$), while the largest measured for the soft abandonment task is r $=0.222$ (p $=0.493$), both of which indicate weak and non-significant correlations. In turn, this strengthens the results reported in Table 5, including those about subreddit diversity, as the variable results in the table cannot be simply explained by the fact that some users participated in multiple subreddits, but rather to their inherent differences. The comparison between the results obtained without LOOCV and the LOOCV aggregate results sheds light on the capacity of our model to generalize to unseen data. To this end, results in Table 5 show a moderate loss in performance in the LOOCV experiment, which is expected given the aforementioned differences among the subreddits. However, the percentage loss is generally contained. For example, in the hard abandonment task the average F1 score on the positive class obtained with the LOOCV is 60.0% of the original one. The loss for the overall metrics is even lower, with the average micro F1 reaching 85.0% of the original value. Similar percentage losses are obtained for the soft abandonment task in the overall metrics, with the LOOCV average positive F1 maintaining 90.7% of the original value. Overall, these figures suggest that the model maintains encouraging performance even when applied to unseen data.
(a) Task: hard abandonment. Feature importance: SHAP.
(b) Task: soft abandonment. Feature importance: SHAP.
(c) Task: hard abandonment. Feature importance: PI.
(d) Task: soft abandonment. Feature importance: PI.